40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 3 18:42:43 2014, Jenne, Update, LSC, Green PDH box boosts BoostCartoon.pdf
    Reply  Thu Sep 4 00:56:44 2014, Jenne, Update, LSC, Green PDH box boosts TFstage_newDesign_3Sept2014.pngBoostNoBoost_newDesign_3Sept2014.pngBoostOnly_newDesign_3Sept2014.pngD0901351-v1_3Sept2014.pdf
       Reply  Thu Sep 4 03:12:55 2014, ericq, Update, LSC, Green PDH box boosts boostedTF.pdfgain.jpg
       Reply  Thu Sep 4 16:45:10 2014, Jenne, Update, LSC, Green PDH box boosts IMG_1712.JPGIMG_1714.JPG
          Reply  Fri Sep 5 05:32:57 2014, Jenne, Update, LSC, Green PDH box boosts TRX_TRY_ALSvsIR_4Sept2014.pdf
             Reply  Fri Sep 5 21:13:57 2014, Jenne, Update, LSC, Green PDH out of loop ALS_XY_outOfLoop_5Sept2014.pdf
Message ID: 10458     Entry time: Fri Sep 5 05:32:57 2014     In reply to: 10452     Reply to this: 10462
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Green PDH box boosts 

[Rana, Jenne, EricQ]

* Too much gain overall on Yend box, needed attenuator on output to get lock.  Rethought gain allocation.  Resoldered board, installed, Ygreen locks nicely.  Error point and control point spectra, box TF and open loop TF data collected, to be plotted.

* Q replaced the Xend box, with a matching TF.

* Locked both arms individually, Yend has lots of low freq fluctuation, Xend has some.  Can't do out of loop measurement since we're going well beyond the range of the PDH signals (Yarm RIN is between 1/2 and 1.) Plot TRX and TRY spectra with ALS lock vs. IR lock to get an idea of what frequencies we have a problem with.

* Tried comm/diff locking anyway.  Works.  Used cm_up script to get CARM to sqrtInvTrans.  Went to powers of about 0.5 (hard to say really, because of fluctuations), put sine at 611.1 Hz, 200 cts onto ETMs (-1*x, +1*y), looked at TF between ALS diff and AS55Q.  Put that amount into the static power normalization spot for AS55.  In steps of 0.1, reduced ALSdiff input matrix elements and increased AS55->DARM element.  2 (3?) times was able to get to AS55Q for DARM.  Lost lock once unknown reason, while reducing CARM offset.  Lost lock once trying to turn on FM4 LSC boost for DARM.

TRX/TRY spectra:

TRX_TRY_ALSvsIR_4Sept2014.pdf

ELOG V3.1.3-